Question Library
Browse, filter, and study verified past examination questions with step-by-step mathematical reasoning.
Searching 18,000+ Questions...
Sifting topics, KaTeX formulas & verified solutions
Showing 5941–5955 of 29821 verified questions
Sort by:
Physics
WAEC / WASSCE 2018
Reflection, Refraction & Dispersion of Light
Diagram
Solution
You are provided with a metre rule, lens, screen, ray box, and other necessary apparatus.
i. Set up the experiment as shown in the diagram above. Measure and record the diameter a\(_{0}\), of the illuminated object.
ii. Place the object at a distance x= 25cm from the lens. Adjust the screen until a sharp image is obtained on the screen.
iii. Measure and record the diameter, a, of the image.
iv. Measure and record the distance v between the lens and the screen.
v. Evaluate y = P = \(\frac{1+y^{2}}{y}\) and T= x+v.
vi. Repeat the procedure for x = 30cm, 35cm, 40cm and 45cm. In each case, determine the corresponding values of a,v,y, P and T.
vii. Tabulate your results.
viii. Plot a graph of P on the vertical axis against T on the horizontal axis starting both axes from the origin (0,0).
ix. Determine the slope, s, of the graph.
x. Determine the intercept, c, on the horizontal axis.
xi Evaluate K = \(\frac{c}{2}\)
xii. State two precautions taken to ensure accurate results.
(b)i. Explain the statement, the focal length of a converging lens is 20cm.
ii. An object is placed at a distance x from a converging lens of focal length 20cm. If the magnification of the real image is 5, calculate the value of x.
Attached Figure / Diagram
medium Difficulty
Verified Correct Answer: Option
Physical Principles & Numerical Solution
Table of value/observation
\(a_0\) = 2.00cm
S/N
x(cm)
a(cm)
v(cm)
y=\(\frac{a}{a_{0}}\)
p=\(\frac{1+y^{2}}{y}\)
T=x+v(cm)
1
25.0
2.60
49.0
1.30
2.07
74.00
2
30.0
2.40
46.0
1.20
2.03
76.00
3
35.0
2.10
45.0
1.05
2.01
80.00
4
40.0
1.80
43.0
0.90
2.01
83.00
5
45.0
1.60
40.0
0.80
2.05
85.00
viii.
viii. Slopes(s) = \(\frac{x_{2} - x_{1}}{y_{2} - y_{1}}\) = \(\frac{76 - 70}{2.073 - 2.040}\)
= \(\frac{6}{0.033}\) = 181.8
x. Intercept (c) = 81
xi. Evaluation k = \(\frac{c}{2}\)
= \(\frac{81}{2}\)
= 40.5
xii. Precautions;
(a) Conxial arrangement of optical instruments
(b) Avoided parallax error in reading metre rule
(c) Lens kept upright
(d) Repeated readings shown on the table
(e) Surface of lens cleaned
(f) Avoided zero error on metre rule
(b)i. The distance between the optical centre and the principal focus of the lens is 20cm or if the parallel beam of light is incident on the lens, it is brought to a focus at a point 20cm from the lens.
ii. M = 5cm, F = 20c, x = u
M = \(\frac{v}{u}\)
5 = \(\frac{v}{u}\)
v = 5u
Using the formula
\(\frac{1}{v} + \frac{1}{u} = \frac{1}{f}\)
\(\frac{1}{5u} + \frac{1}{u} = \frac{1}{20}\)
\(\frac{1+5}{5u} = \frac{1}{20}\)
5u = 20(1 + 5)
5u = 20 x 6
5u = 120
u = \(\frac{120}{5}\) = 24
x = 24cm
Source: WAEC / JAMB / NECO Classroom Archive (Record #59884)
Chemistry
UTME / JAMB 2018
Kinetic Theory of Matter & Gas Laws
Solution
A given amount of gas occupies 10.0dm\(^3\) at 4 atm and 273°C. The number of moles of the gas present is?
[Molar volume of gas at s.t.p = 22.4 dm\(^3\)]
A
0.89 mol
B
1.90 mol
C
3.80 mol
D
5.70 mol
medium Difficulty
Verified Correct Answer: Option A
Chemical Principles & Mechanism
Given:
P = 4 atm , V = 10 dm\(^3\) , n = ? R = 0.082 (constant) T = 273 + 273 = 546K
Using the ideal gas law, PV = nRT
n = \(\frac{PV}{RT}\)
n = \(\frac{{4}\times{10}}{{0.082}\times{546}}\)
n = \(\frac{40}{44.772}\)
n = 0.893 mole
Source: WAEC / JAMB / NECO Classroom Archive (Record #47284)
Chemistry
UTME / JAMB 2018
Water, Solutions & Solubility
Solution
Which of these sources of water may likely contain the least concentration of Ca\(^{2+}\) and Mg\(^{2+}\)?
A
Spring water
B
River water
C
Tap water
D
Sea water
medium Difficulty
Verified Correct Answer: Option C
Chemical Principles & Mechanism
Tap water is likely to contain less Ca\(^{2+}\) and Mg\(^{2+}\). The reason is because looking through all the given options, spring water, river water and sea water are natural water that has not undergone any treatment. Conversely, it is believed that before tap water is distributed for consumption, it must have been treated, thereby reducing the Ca\(^{2+}\) and Mg\(^{2+}\) content in the water.
Source: WAEC / JAMB / NECO Classroom Archive (Record #47654)
Mathematics
UTME / JAMB 2018
Sets, Logic & Venn Diagrams
Solution
In a class of 40 students, 32 offer Mathematics, 24 offer Physics and 4 offer neither Mathematics nor Physics. How many offer both Mathematics and Physics?
A
4
B
8
C
16
D
20
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
Let the number of people that offer both Mathematics and Physics = y
Then, \((32 - y) + y + (24 - y) + 4 = 40\)
\(60 - y = 40 \implies y = 20\)
\(\therefore\) 20 students offer both Mathematics and Physics.
Source: Myschool Past Questions Archive (Record #46974)
Mathematics
UTME / JAMB 2018
Plane Geometry, Polygons & Circle Theorems
Solution
Find the values of x for which
\(\frac {x+2}{4}\) - \(\frac{2x - 3}{3}\) < 4
A
x < 8
B
x > -6
C
x < 4
D
x > -3
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
\(\frac {x+2}{4}\) - \(\frac{2x - 3}{3} < 4\)
\(\frac{3(x + 2) - 4(2x - 3)}{12} < 4\)
\(3x + 6 - 8x + 12 < 48 \)
\(18 - 5x < 48 \implies -5x < 30\)
\(\therefore x > -6\)
Source: Myschool Past Questions Archive (Record #46975)
Mathematics
UTME / JAMB 2018
Statistics (Data Representation, Central Tendency & Dispersion)
Diagram
Solution
The pie chart shows the monthly expenditure of a public servant. The monthly expenditure on housing is twice that of school fees. How much does the worker spend on housing if his monthly income is N7200?
Attached Figure / Diagram
A
1000
B
2000
C
3000
D
4000
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
Let the angle for school fees = x^\circ
Then Housing = 2x^\circ
120^\circ + 90^\circ + x^\circ + 2x^\circ = 360^\circ
3x^\circ = 150^\circ \(\implies\) x^\circ = 50^\circ.
Amount spent on housing = \(\frac{100}{360} \times 7200\)
= N2000.
Source: Myschool Past Questions Archive (Record #46978)
Mathematics
UTME / JAMB 2018
Fractions, Decimals, Percentages & Approximations
Solution
A trader realises 10x - x\(^2\) Naira profit from the sale of x bags of corn. How many bags will give him the maximum profit?
A
7
B
6
C
5
D
4
medium Difficulty
Verified Correct Answer: Option C
Step-by-Step Mathematical Solution
Profit (P) = 10\(_x\) − \(_x\)2
Maximum profit can be achieved when the differential of profit with respect to number of bags(x) is 0
i.e. \(\frac{dp}{dx}\) = 0
\(\frac{dp}{dx}\) = 10 - 2x = 0
10 = 2x
Then x = \(\frac{10}{2}\) = 5
Answer is C
Source: Myschool Past Questions Archive (Record #46980)
Mathematics
UTME / JAMB 2018
Number Bases & Modular Arithmetic
Solution
If y = 23\(_{five}\) + 101\(_{three}\) , find y, leaving your answer in base two
A
1110
B
10111
C
11101
D
111100
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
y = 23\(_{five}\) + 101\(_{three}\)
23\(_{five}\) = \(2 \times 5^1 + 3 \times 5^0\)
= 13\(_{ten}\)
101\(_{three}\) = \(1 \times 3^2 + 0 \times 3^1 + 1 \times 3^0\)
= 10\(_{ten}\)
y\(_{ten}\) = 13\(_{ten}\) + 10\(_{ten}\)
= 23\(_{ten}\)
= 10111\(_{two}\)
Source: Myschool Past Questions Archive (Record #46982)
Mathematics
UTME / JAMB 2018
Plane Geometry, Polygons & Circle Theorems
Diagram
Solution
Find the value of x in the diagram
Attached Figure / Diagram
A
10°
B
28°
C
36°
D
40°
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
The diagram shows angles at a point, the total angle at a point is 360
x - 10 + 4x - 50 + 2x + 3x + 20 = 360
10x - 40 = 360
10x = 360 + 40
10x = 400
x = \(\frac{400}{10}\)
x = 40
Source: Myschool Past Questions Archive (Record #46983)
Mathematics
UTME / JAMB 2018
Plane Geometry, Polygons & Circle Theorems
Solution
Solve for t in the equation \(\frac{3}{4}\)t + \(\frac{1}{3}\)(21 - t) = 11
A
\(\frac{9}{13}\)
B
\(\frac{7}{13}\)
C
5
D
9\(\frac{3}{5}\)
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
\(\frac{3}{4}\) t + \(\frac{1}{3}\) (21 - t) = 11
Multiply through by the LCM of 4 and 3 which is 12
12 x(\(\frac{3}{4}\) t) + 12 x (\(\frac{1}{3}\) (21 - t)) = (11 x 12)
9t + 4(21 - t) = 132
9t + 84 - 4t = 132
5t + 84 = 132
5t = 132 - 84 = 48
t = \(\frac{48}{5}\)
t = 9 \(\frac{3}{5}\)
Answer is D
Source: Myschool Past Questions Archive (Record #46984)
Mathematics
UTME / JAMB 2018
Plane Geometry, Polygons & Circle Theorems
Solution
A school girl spends \(\frac{1}{4}\) of her pocket money on books and \(\frac{1}{3}\) on dress. What fraction remains?
A
\(\frac{5}{6}\)
B
\(\frac{7}{12}\)
C
\(\frac{5}{12}\)
D
\(\frac{1}{6}\)
medium Difficulty
Verified Correct Answer: Option C
Step-by-Step Mathematical Solution
let the girls' initial pocket money be whole(1)
\(\frac{1}{4}\) of 1 = books, \(\frac{1}{3}\) of 1 = dress
fraction of her money left = 1 - \(\frac{1}{4}\) - \(\frac{1}{3}\)
= \(\frac{5}{12}\)
Source: Myschool Past Questions Archive (Record #46985)
Mathematics
UTME / JAMB 2018
Algebraic Expressions, Polynomials & Factorisation
If \(\frac{x}{a + 1}\) + \(\frac{y}{b}\) = 1. Make y the subject of the relation.
A
\(\frac{b(a + 1 - x)}{a + 1}\)
B
\(\frac{a + 1}{b(a - x + 1)}\)
C
\(\frac{a(b - x + 1)}{b + 1}\)
D
\(\frac{b}{a(b - x + 1)}\)
medium Difficulty
Verified Correct Answer: Option A
Step-by-Step Mathematical Solution
No expanded working has been attached for this question yet. The verified answer is Option A.
Source: Myschool Past Questions Archive (Record #46987)
Mathematics
UTME / JAMB 2018
Mensuration (Areas, Perimeters & Volumes)
Solution
Calculate the total surface area of a cupboard which measures 12cm by 10cm by 8cm
A
1920cm\(^2\)
B
592cm\(^2\)
C
296cm\(^2\)
D
148cm\(^2\)
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
Total surface area of a cupboard is given by the equation A = 2(lb + bh + lh) L = 12, b = 10, h = 8
A = 2((12 x 10) + (10 x 8) + (12 x 8))
A = 2(120 + 80 + 96)
A = 2 x 296
A = 592cm\(^2\)
Answer is B
Source: Myschool Past Questions Archive (Record #46988)
Mathematics
UTME / JAMB 2018
Fractions, Decimals, Percentages & Approximations
Solution
Convert 0.04945 to two significant figures
A
0.040
B
0.049
C
0.050
D
0.49
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
0.04945 to 2s.f is 0.049
Source: Myschool Past Questions Archive (Record #46989)
Mathematics
UTME / JAMB 2018
Permutations, Combinations & Probability
Solution
The probabilities that John and James pass an examination are \(\frac {3}{4}\)and \(\frac {3}{5}\) respectively. Find the probability of both boys failing the examination.
A
\(\frac {1}{10}\)
B
\(\frac {2}{10}\)
C
\(\frac {9}{20}\)
D
\(\frac {11}{20}\)
medium Difficulty
Verified Correct Answer: Option A
Step-by-Step Mathematical Solution
To find the probability of both John and James failing the examination:
Probabilities of failing:
\(P(\text{John fails}) = 1 - P(\text{John passes}) = 1 - \frac{3}{4} = \frac{1}{4}\)
\(P(\text{James fails}) = 1 - P(\text{James passes}) = 1 - \frac{3}{5} = \frac{2}{5}\)
Probability of both failing:
\(P(\text{Both fail}) = P(\text{John fails}) \times P(\text{James fails}) = \frac{1}{4} \times \frac{2}{5} = \frac{2}{20} = \frac{1}{10}\)
Thus, the probability of both boys failing the examination is:
\(\frac{1}{10}\)
Source: Myschool Past Questions Archive (Record #46991)